# Cross-Modal Learning for Anomaly Detection in Fused Magnesium Smelting Process: Methodology and Benchmark

### v1

@@ V Experiments
In this section, we evaluate the proposed FmFormer on the fused magnesium smelting process benchmark and compare it with several state-of-the-art learning-based anomaly detection methods that apply both unimodal inputs (current or visual) and cross-modal inputs. In the experiments, quantitative evaluations and visual comparisons specifically for dense prediction are performed. In addition, we empirically investigate the modules in the proposed FmFormer through several ablation studies.
V-A Evaluation Metrics
We employ the following performance metrics that are commonly used in anomaly detection:
  • Accuracy ($Acc$) indicates the ratio of correctly predicted samples with respect to all the testing samples, i.e, $Acc=(TP+TN)/(TP+TN+FP+FN)$, where $TP$, $TN$, $FP$ and $FN$ means true positive, true negative, false positive and false negative, respectively.
  • F1-score ($F1$) is a comprehensive metric that considers precision $P=TP/(TP+FP)$ and recall, $R=TP/(TP+FN)$. Then the F1-score is computed as $F1=2P\cdot R/(P+R)$.
  • False Detection Rate (FDR), also known as false positive rate, is defined as the ratio of incorrectly predicted normal samples with respect to all the real normal (negative) samples, i.e., $FDR=FP/(TN+FP)$.
  • Miss Detection Rate (MDR) is also known as false negative rate that indicates the ratio of incorrectly predicted abnormal samples to all the real abnormal (positive) samples, i.e., $MDR=FN/(TP+FN)$.
  • Mean Intersection over Union (mIoU) is employed to evaluate the performance of dense prediction, which is typically used for image segmentation. It is defined as the ratio of the intersection to the union of the pixel-level ground truth and the dense prediction result, i.e., $mIoU=\frac{1}{K}\sum_{k\in K}TP/(TP+FN+FP)$.
TABLE II: Quantitative comparison with state-of-the-art methods on the fused magnesium smelting process benchmark.
@@ V-A Evaluation Metrics
FmFormer (Ours) 0.9836 0.9846 0.0216 0.0116 0.8409
V-B Comparison With State-of-the-Art Methods
We demonstrate the effectiveness of the proposed FmFormer by comparing it with several baseline methods using unimodal settings (current/visual) and a cross-modal setting.
V-B1 Current modality
@@
Table II presents the quantitative comparison with the baseline methods on the proposed fused magnesium smelting process benchmark. For current modal input, the proposed FmFormer achieves comparable performance with state-of-the-art Transformer-based methods that specifically designed for time-series prediction. Nevertheless, from the performance perspective, anomaly detection using pure current information is still far from practical industrial applications.
V-B2 Visual modality
@@
Four state-of-the-art methods or backbones are evaluated, which are 3DCRNN [11], ViT [16], ViVit [28] and TubeVit [29]. 3DCRNN [11] is a typical convolution recurrent-based anomaly detection framework specifically for fused magnesium smelting processes, utilizing 2D convolutional LSTM cells [13] to extract 3D (2D spatial and 1D temporal) features. ViT [16], ViViT [28] and TubeViT [29] are three Transformer-based methods specifically designed for video input, and employ non-overlapping 2D patches [16], 3D tubes [28] and 3D tubes of different shapes [29] for tokenization, respectively.
@@
As shown in Table II, the visual models exploit visual features that are more stable and prominent than current information, which provide them a distinct advantage. Besides, 3DCRNN [11] achieves comparable performance to the Transformer-based methods, ViT [16], ViVit [28] and TubeVit [29], due to the effective spatial-temporal modeling of convolutional LSTM units. Among the unimodality-based methods using visual input, our FmFormer achieves the best comprehensive performance.
@@
Figure 5: Visual comparison of the proposed FmFormer (cross-modality) with three state-of-the-art Transformer-based methods for pixel-level anomaly detection on three challenging cases. In each case, two close frames in a video are displayed to demonstrate the abnormal dynamics. We superimpose the pixel-level detection result of each method with the corresponding frame for better viewing. In the first case, intense flame light of the furnace (first frame) interferes with most detection methods. The anomaly is not detected until the light interference subsided (second frame). In the second case, visual occlusion from heavy water mist affects the compared methods, resulting in miss detection of the subtle anomaly (second frame). In the third case, fluctuations in ambient light influence the accuracy of anomaly location of the compared methods. In these challenging cases, the proposed FmFormer considers current information as a prompt for normal or abnormal conditions, thus demonstrating better robustness under disturbances and temporal consistency under occlusions.
TABLE III: Performance of the FmFormer without and with the dilated tokenization mechanism.
Modality Tokenization Classification Dense prediction
$Acc\uparrow$ $F1\uparrow$ $FDR\downarrow$ $MDR\downarrow$ $mIoU(\%)\uparrow$
Visual w/o dilated tokenization 0.9694 0.9709 0.0217 0.0382 0.8369
w dilated tokenization 0.9731 0.9744 0.0224 0.0307 0.8382
Cross-modal w/o dilated tokenization 0.9757 0.9772 0.0321 0.0171 0.8376
w dilated tokenization 0.9836 0.9846 0.0216 0.0116 0.8409
V-B3 Cross-modality
@@
Six state-of-the-art methods are evaluated, including DCNN-SVM [8], SSFGGAN [45], Unicoder-VL [64], ClipBERT [47], CAPTURE [65] and BiLM [49]. DCNN-SVM [8] and SSFGGAN [45] achieve multi-modal learning through linear fusion of features from different modalities, which are designed specifically for anomaly detection in fused magnesium smelting processes. Unicoder-VL [64] encodes vectorized 2D image patches conjointly with 1D sequence (current) into Transformer backbones. ClipBERT [47] and BiLM [49] first employ pre-trained backbones to extract visual features from the video modality, and then embed the visual features together with 1D sequence into Transformers. The different is that ClipBERT [47] squeezes the temporal dimension of visual features via an average-pooling for efficiency. CAPTURE [65] also adopts a pipeline of self-attention and cross-attention for cross-modal learning. The main difference between our FmFormer and CAPTURE [65] is that we use a cascading structure to alternatively stack self-attention and cross-attention. This structure helps the network to capture internal features in each modality and correlation features across modalities in a progressive manner. Although most of the baseline methods are designed for classification task, we extend the Transformer-based models to dense prediction task by assembling tokens into image-like features as introduced in Section III-C1.
@@
As indicated by the quantitative results in Table II, the Transformer-based models using cross-modal input (Unicoder-VL [64], ClipBERT [47] and BiLM [49]) generally outperform models using single visual modality (ViT [16], ViVit [28] and TubeVit [29]). Compared with conventional CNN-based methods using linear fusion (DCNN-SVM [8] and SSFGGAN [45]), Transformer-based methods (Unicoder-VL [64], ClipBERT [47] and BiLM [49]) generally demonstrates higher performance, as shown in Table II since the explicit correlation modeling achieved via Transformer encoder. Among methods using cross-modal input, the proposed FmFormer achieves superior performances, confirming its effectiveness for both class-level prediction and pixel-level prediction.
V-B4 Dense prediction
@@
Fig. 5 visualizes the qualitative comparison of dense predictions of the proposed FmFormer with three state-of-the-art Transformer-based methods, ViViT [28] (visual), ClipBERT [47] (cross-modal), and BiLM [49] (cross-modal), on three representative cases. In the first case, intense flame light of the furnace has stronger visual saliency than the anomaly region (please see the first frame). The compared methods fail to detect the anomaly under the strong disturbance until the disturbance subsides (second frame). In the second case, heavy water mist occludes the visual features of the anomaly, so the detection method can only rely on the information from the previous frames and three-phase alternating current. In this case, ViViT [28] with only visual input fails to detect the anomaly. In the first frame, ClipBERT [47] and BiLM [49] with cross-modal input detect the anomaly successfully, but predict the wrong location. In the second frame, they also fail to spot the abnormal condition. In the third case, fluctuations of ambient light lead to variations of visual features, which in turn cause the compared methods to misestimate the anomaly regions (see ClipBERT [47] and BiLM [49]) or even produce missed detection (see ViViT [28]). In comparison, the proposed FmFormer leverages current information as a prompt for normal or abnormal conditions, thereby demonstrating robustness under disturbances and temporal consistency under occlusions in complex industrial environment.
V-C Ablation Study
Figure 6: Performance of the FmFormer under different settings of input sequence lengths (horizontal coordinate).
V-C1 Sequence length
Benefiting from the designed structure, our FmFormer can be adapted to input of variable sequence length without network retraining or fine-tuning [^1] Since we employ learnable positional embeddings for visual input and current input, bilinear interpolation is used to accommodate inputs of different sequence lengths.. It should be noted that since the input sequences of video and current represent the operating state at the current instant and how that state may change, we only need to ensure the final frame and three-phase current values in the input sequences belong to the same instant, while the sequence lengths of the two modalities do not need to be exactly the same. The length of the input sequence of each modality influences the network to perceive the dynamic features of normal or abnormal conditions.
In this experiment, we investigate the performance of our FmFormer under different settings of input sequence lengths, as shown in Fig 6. For the visual modality, the performance of the proposed method continues to increase when the video sequence length is longer than 1.5 seconds. However, the MDR tends to increase when the input sequence is too long. So we chose an input sequence length of 1.5 seconds. For the current modality, the performance of the proposed method improves as the sequence length increases since it is less influenced by the current noise. However, excessively long sequences of current input may drive the self-attention mechanism to tend to focus on the working conditions of other time series instead of the current moment, resulting in the performance degradation. In this experiment, the best performance is achieved when the length of the input current sequence is around 220 seconds.
V-C2 Multiscale tokenization
In this experiment, the effectiveness of the proposed multiscale tokenization module is studied by comparing the results without and with the dilated tokenization mechanism. Note that the multiscale feature blending formulated in (10) is also removed when the dilated tokenization mechanism is not utilized. As shown in Table III, the models using the proposed multiscale tokenization module achieve overall better performance in both single (visual) modality and cross-modal input settings, which demonstrate the effectiveness of the proposed module.
@@ V-C2 Multiscale tokenization
TABLE IV: Performance of the FmFormer using unimodal and cross-modal input, where “LF” indicates the simple linear fusion mechanism in (12).
@@
V-C3 Unimodality vs. cross-modality
@@
To investigate the influence of each modality on the performance of anomaly detection, we degrade the proposed FmFormer to a unimodal method by using only one modality as input. Due to the absence of cross-modal input, we replace the MHCA with the MHSA in the Transformer encoder and remove the linear fusion mechanism in (12). Table IV lists the performance of the proposed FmFormer using a single visual or current modality. As expected, the visual model outperforms the current model because the visual features are more stable than the current features.
TABLE V: Performance of the FmFormer under different settings of modality interaction mechanisms. The first row (ID 0) serves as a baseline using the single visual modality. Since the linear fusion mechanism does not interact with the dense prediction head, the dense prediction results with/without it are the same.
ID Input Mechanism Classification Dense prediction
MHSA MHCA Linear fusion $Acc\uparrow$ $F1\uparrow$ $FDR\downarrow$ $MDR\downarrow$ $mIoU(\%)\uparrow$
0 Visual 0.9731 0.9744 0.0224 0.0307 0.8382
1 Cross-modality 0.9748 0.9764 0.0354 0.0159 0.8382
2 Cross-modality 0.9711 0.9726 0.0253 0.0318 0.7627
3 Cross-modality 0.9757 0.9769 0.0187 0.0291 0.8409
4 Cross-modality 0.9836 0.9846 0.0216 0.0116 0.8409
@@
In this study, the performance of adopting cross-modal input is also verified by simply integrating the two models with unimodal inputs through a linear fusion mechanism in (12). Note that no further training or fine-tuning is performed for the linear fusion. As shown in Table IV, despite the poor performance of the current model, it can be integrated with the visual model to improve the comprehensive performance, especially the MDR decreased by nearly half compared to the vision model. This ablation study fully verifies that cross-modal input can effectively enhance anomaly detection.
V-C4 Effectiveness of modality interaction mechanisms
In this experiment, we validate the effectiveness of the applied modality interaction mechanisms, including the MHSA for internal features encoding within each modality, the MHCA for correlation feature exploration across modalities, and the simple linear fusion mechanism for explicitly weighting the predictions from two modalities. For the settings with only MHSA or MHCA, we replicate the specific interaction mechanism to keep the numbers of the model parameters the same.
Table V lists the performance of the proposed FmFormer using different settings of modality interaction mechanisms, where the first row (ID 0) serves as a baseline using the single visual modality. It should be noted that the linear fusion mechanism does not interact with the dense prediction head, and thus, the dense prediction results with/without it are the same (e.g., ID 0 vs. ID 1 and ID 3 vs. ID 4). As mentioned in Section V-C3, utilizing a simple linear fusion of the two modalities (ID 1) is also able to improve the model performance compared to the single modality (ID 0). While in comparison, the model with only the MHCA (ID 2) has an unsatisfactory performance and is even worse than the baseline setting using a single modal input (ID 0), especially for dense prediction tasks. It indicates that directly modeling the cross-modal correlation by using cross-attention without self-attention feature encoding degrades the model performance. The self-attention provides sufficient preparation for correlation exploration in the cross-attention. In addition, the model with the MHCA (ID 3) outperforms the model with the linear fusion mechanism (ID 1), which verifies the effectiveness of this explicit modality interaction mechanism. Among these settings, the FmFormer with the full modality interaction mechanisms (ID 4) achieves the best comprehensive performance.
VI Conclusion and Limitations

### v2

@@ V Experiments
In this section, we evaluate the proposed FmFormer (principally the base model) on the fused magnesium smelting process benchmark and compare it with several state-of-the-art learning-based anomaly detection methods that apply both unimodal inputs (current or visual) and cross-modal inputs. In the experiments, quantitative evaluations and visual comparisons specifically for dense prediction are performed. In addition, we empirically investigate the modules in the proposed FmFormer through several ablation studies. More applications to general industrial anomaly detection are explored in the supplementary material.
We employ four commonly used metrics [58], accuracy, F1-score (F1), False Detection Rate (FDR), and Miss Detection Rate (MDR), for class-level anomaly detection. For pixel-level anomaly detection, we use Mean Intersection over Union (mIoU) [59]. False Detection Rate (FDR), also known as false positive rate, is defined as the ratio of incorrectly predicted normal samples with respect to all the real normal (negative) samples. Miss Detection Rate (MDR) is also known as false negative rate that indicates the ratio of incorrectly predicted abnormal samples to all the real abnormal (positive) samples.
TABLE III: Quantitative comparison with state-of-the-art methods on the fused magnesium smelting process benchmark.
@@
FmFormer-B (Ours) 0.9837 0.9847 0.0217 0.0117 0.8409
  • 1
    The best results are highlighted in bold, and the second-best results are underlined.
V-A Comparison With State-of-the-Art Methods
We evaluate the effectiveness of our base model, FmFormer-B, by comparing it with several baseline methods using unimodal settings (current/visual) and a cross-modal setting.
V-A1 Current modality
@@
Table III presents the quantitative comparison with the baseline methods on the proposed fused magnesium smelting process benchmark. For current modal input, the proposed FmFormer-B achieves comparable performance with state-of-the-art Transformer-based methods that specifically designed for time-series prediction. Nevertheless, from the performance perspective, anomaly detection using pure current information is still far from practical industrial applications.
V-A2 Visual modality
@@
Four state-of-the-art methods or backbones are evaluated, which are 3DCRNN [12], ViT [16], ViViT [28] and TubeVit [29]. 3DCRNN [12] is a typical convolution recurrent-based anomaly detection framework specifically for fused magnesium smelting processes, utilizing a 2D convolutional LSTM to extract 3D (2D spatial and 1D temporal) features. ViT [16], ViViT [28] and TubeViT [29] are three Transformer-based methods specifically designed for video input, and employ non-overlapping 2D patches [16], 3D tubes [28] and 3D tubes of different shapes [29] for tokenization, respectively.
@@
As shown in Table III, the visual models exploit visual features that are more stable and prominent than current information, which provide them a distinct advantage. Besides, 3DCRNN [12] achieves comparable performance to the Transformer-based methods, ViT [16], ViViT [28] and TubeVit [29], due to the effective spatial-temporal modeling of convolutional LSTM units. Among the unimodality-based methods using visual input, our FmFormer-B achieves the best comprehensive performance.
@@
Figure 5: Visual comparison of the proposed FmFormer-B (cross-modality) with three state-of-the-art Transformer-based methods for pixel-level anomaly detection on three challenging cases. In each case, two close frames in a video are displayed to demonstrate the abnormal dynamics. We superimpose the pixel-level detection result of each method with the corresponding frame for better viewing. In the first case, intense flame light of the furnace (first frame) interferes with most detection methods. The anomaly is not detected until the light interference subsided (second frame). In the second case, visual occlusion from heavy water mist affects the compared methods, resulting in miss detection of the subtle anomaly (second frame). In the third case, fluctuations in ambient light influence the accuracy of anomaly location of the compared methods. In these challenging cases, the proposed FmFormer-B considers current information as a prompt for normal or abnormal conditions, thus demonstrating better robustness under disturbances and temporal consistency under occlusions.
V-A3 Cross-modality
@@
Six state-of-the-art methods are evaluated, including DCNN-SVM [8], SSFGGAN [44], Unicoder-VL [64], ClipBERT [46], CAPTURE [65] and BiLM [47]. DCNN-SVM [8] and SSFGGAN [44] achieve multi-modal learning through linear fusion of features from different modalities, which are designed specifically for anomaly detection in fused magnesium smelting processes. Unicoder-VL [64] encodes vectorized 2D image patches conjointly with 1D sequence (current) into Transformer backbones. ClipBERT [46] and BiLM [47] first employ pre-trained backbones to extract visual features from the video modality, and then embed the visual features together with 1D sequence into Transformers. The different is that ClipBERT [46] squeezes the temporal dimension of visual features via an average-pooling for efficiency. CAPTURE [65] also adopts a pipeline of self-attention and cross-attention for cross-modal learning. Since most baseline methods are designed for classification task, we extend these Transformer-based models to dense prediction task by assembling tokens into image-like features as introduced in Section III-C1.
@@
As indicated by the quantitative results in Table III, the Transformer-based models using cross-modal input (Unicoder-VL [64], ClipBERT [46] and BiLM [47]) generally outperform models using single visual modality (ViT [16], ViViT [28] and TubeVit [29]). Compared with conventional CNN-based methods using linear fusion (DCNN-SVM [8] and SSFGGAN [44]), Transformer-based methods (Unicoder-VL [64], ClipBERT [46] and BiLM [47]) generally demonstrates higher performance, as shown in Table III since the explicit correlation modeling achieved via Transformer encoder. Among methods using cross-modal input, the proposed FmFormer-B achieves superior performances, confirming its effectiveness for both class-level prediction and pixel-level prediction. Moreover, the results indicate that while the video modality provides significant anomaly features, the current modality is crucial for enhancing anomaly detection in the fused magnesium smelting process.
V-A4 Dense prediction
@@
Fig. 5 visualizes the qualitative comparison of dense predictions of the proposed FmFormer-B with three state-of-the-art Transformer-based methods, ViViT [28] (visual), ClipBERT [46] (cross-modal), and BiLM [47] (cross-modal), on three representative cases. In the first case, intense flame light of the furnace has stronger visual saliency than the anomaly region (please see the first frame). The compared methods fail to detect the anomaly under the strong disturbance until the disturbance subsides (second frame). In the second case, heavy water mist occludes the visual features of the anomaly, so the detection method can only rely on the information from the previous frames and three-phase alternating current. In this case, ViViT [28] with only visual input fails to detect the anomaly. In the first frame, ClipBERT [46] and BiLM [47] with cross-modal input detect the anomaly successfully, but predict the wrong location. In the second frame, they also fail to spot the abnormal condition. In the third case, fluctuations of ambient light lead to variations of visual features, which in turn cause the compared methods to misestimate the anomaly regions (see ClipBERT [46] and BiLM [47]) or even produce missed detection (see ViViT [28]). In comparison, the proposed FmFormer-B leverages current information as a prompt for normal or abnormal conditions, thereby demonstrating robustness under disturbances and temporal consistency under occlusions in complex industrial environment.
@@
TABLE IV: Performance of the FmFormer-B using cross-modal and unimodal input, where “LF” denotes the linear fusion in (12).
@@
TABLE V: Performance with respect to model capacity (configuration).
$Acc\uparrow$ Param. # FLOPs FPS $\uparrow$
FmFormer-T 0.9676 0.28M 110M 91.5
FmFormer-S 0.9778 0.48M 174M 91.2
FmFormer-B 0.9837 1.86M 610M 80.1
FmFormer-L 0.9839 3.64M 1,205M 48.7
V-B Ablation Study
V-B1 Unimodality vs. cross-modality
@@ V-B Ablation Study
To investigate the influence of each modality on the performance of anomaly detection, we degrade the proposed FmFormer to a unimodal method by using only one modality as input. Due to the absence of cross-modal input, we replace the MHCA with the MHSA in the Transformer encoder and remove the LF mechanism in (12). Table IV lists the performance of the proposed FmFormer-B using a single visual or current modality. As expected, the visual model outperforms the current model because the visual features are more stable than the current features.
@@
In this study, the performance of adopting cross-modal input is also verified by simply integrating the two models with unimodal inputs through a LF mechanism in (12). Note that no further training or fine-tuning is performed for the LF mechanism. As shown in Table IV, despite the poor performance of the current model, it can be integrated with the visual model to improve the comprehensive performance, especially the MDR decreased by nearly half compared to the vision model. This ablation study fully verifies that cross-modal input can effectively enhance anomaly detection.
V-B2 Performance with respect to model capacity
We investigate the performance of our models under different configurations (model capacities), with detailed specifications listed in Table II. As shown in Table V, the prediction accuracy improves from 0.9676 (FmFormer-T) to 0.9839 (FmFormer-L), demonstrating the scalability of FmFormer. Additionally, the inference latency, tested on an NVIDIA TESLA V100, remains satisfactory, highlighting the potential for industrial deployment.
Figure 6: Performance of the FmFormer-B under different settings of input sequence lengths (horizontal coordinate).
TABLE VI: Performance of the FmFormer-B without and with the dilated tokenization mechanism.
Modality Dilated tokenization $Acc\uparrow$ $mIoU(\%)\uparrow$
Visual 0.9694 0.8369
0.9731 0.8382
Cross-modal 0.9757 0.8376
0.9837 0.8409
V-B3 Sequence length
Benefiting from the designed structure, our FmFormer can be adapted to input of variable sequence length without network retraining or fine-tuning [^1] We use bilinear interpolation for the positional embeddings to accommodate inputs of different sequence lengths.. It should be noted that since the input sequences of video and current represent the operating state at the current instant and how that state may change, we only need to ensure the final frame and three-phase current values in the input sequences belong to the same instant, while the sequence lengths of the two modalities do not need to be exactly the same. The length of the input sequence of each modality influences the network to perceive the dynamic features of normal or abnormal conditions.
TABLE VII: Performance of the FmFormer-B under different settings of the cross-modal Transformer encoder.
ID MHSA MHCA (uni.) MHCA (bi.) $Acc\uparrow$ $mIoU(\%)\uparrow$
0 0.9748 0.8382
1 0.9694 0.7562
2 0.9711 0.7627
3 0.9761 0.8365
4 0.9837 0.8409
In this experiment, we investigate the performance of our FmFormer-B under different settings of input sequence lengths, as shown in Fig 6. For the visual modality, the performance of the proposed method continues to increase when the video sequence length is longer than 1.5 seconds. However, the MDR tends to increase when the input sequence is too long. So we chose an input sequence length of 1.5 seconds. For the current modality, the performance of the proposed method improves as the sequence length increases since it is less influenced by the current noise. However, excessively long sequences of current input may drive the self-attention mechanism to tend to focus on the working conditions of other time series instead of the current moment, resulting in the performance degradation. In this experiment, the best performance is achieved when the length of the input current sequence is around 220 seconds.
V-B4 Effectiveness of multiscale tokenization
In this experiment, the effectiveness of the proposed multiscale tokenization module is studied by comparing the results without and with the dilated tokenization mechanism. Note that the multiscale feature blending formulated in (10) is also removed when the dilated tokenization mechanism is not utilized. As shown in Table VI, the models using the proposed multiscale tokenization module achieve overall better performance in both single (visual) modality and cross-modal input settings, which demonstrate the effectiveness of the proposed module.
V-B5 Effectiveness of cross-modal Transformer encoder
We validate the effectiveness of our encoder, which includes the MHSA for internal features encoding within each modality, the MHCA for correlation feature exploration across modalities. The MHCA mechanism includes a unidirectional form, referred to as “MHCA (uni.)” for short, and a bidirectional form, denoted as “MHCA (bi.).” For configurations using only MHSA or MHCA, we replicate the interaction mechanism to maintain consistent model parameters. Results in Table VII show that the model with only bidirectional MHCA (ID 2) outperforms the unidirectional (current-to-visual) version (ID 1). However, both forms of solely MHCA perform worse than the model using only MHSA (ID 0), particularly in dense prediction tasks. When combining MHSA with bidirectional MHCA (ID 4), we observe higher accuracy in both classification and dense prediction compared to the unidirectional model (ID 3).
These findings indicate that i) the bidirectional MHCA mechanism effectively enhances class-level and pixel-level prediction tasks; and ii) directly modeling the cross-modal correlation by using cross-attention without self-attention feature encoding degrades the model performance. The self-attention provides sufficient preparation for correlation exploration in cross-attention.
V-B6 Effectiveness of multi-head decoder
We also evaluate various decoder designs, including the classification head (abbreviated as “cls. head”) with or without the LF mechanism, as well as the dense prediction head (abbreviated as “dense head”)), as shown in Table VIII. Using only the single dense prediction head (ID 0), our FmFormer-B suffers a performance degradation in mIoU. Additionally, the single classification head (ID 1) performs worse compared to our full model with the multi-head decoder (ID3). These performance losses are essentially due to training with a single loss term ($\mathcal{L}^{pix}$ or $\mathcal{L}^{cls}$). Furthermore, when comparing the model without LF mechanism (ID 2) to the version with it (ID 3), the latter demonstrates significantly better performance. Since the LF mechanism does not interact with the dense prediction head, the mIoU values are the same with or without it. These results highlight the effectiveness of the proposed multi-head decoder.
TABLE VIII: Performance of the FmFormer-B under different settings of the multi-head decoder.
ID Cls. head Dense head LF $Acc\uparrow$ $mIoU(\%)\uparrow$
0 - 0.8276
1 0.9795 -
2 0.9757 0.8409
3 0.9837 0.8409
VI Conclusion and Discussion